OpenStudy (babynini):

Find two numbers whose difference is 60 and whose product is a minimum

OpenStudy (babynini):

@jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (babynini):

em well I tried to set up the equations but I think that's where i'm going off x-y=60 xy= ?

jimthompson5910 (jim_thompson5910):

if `x-y=60` then we can solve for y to get `y = x-60`, agreed?

OpenStudy (babynini):

Wait I think that x-y=60 is incorrect If the difference is 60 then it should be x+60=y

jimthompson5910 (jim_thompson5910):

x+60=y leads to 60 = y-x

jimthompson5910 (jim_thompson5910):

if x is the bigger number then x-y = 60

OpenStudy (babynini):

hm? if the difference is 60 then if we were to do x-y that would give us a smaller number than 60.. if that makes sense. or maybe let's just try this out haha

jimthompson5910 (jim_thompson5910):

`difference is 60` just means we subtract the two values (x & y) to get 60 we could do y-x or x-y, but let's go with x-y

OpenStudy (babynini):

okay

OpenStudy (babynini):

then we plug that into the x(y) function, yeah? f(x)=x(x-60) f'(x)=2x-60

jimthompson5910 (jim_thompson5910):

yes f(x) = x(x-60) f(x) = x^2 - 60x f ' (x) = 2x - 60

OpenStudy (babynini):

so x = 30

jimthompson5910 (jim_thompson5910):

yes after solving f ' (x) = 0

OpenStudy (babynini):

then plugging that x into the original we get that y=-30 haha that makes so much sense.

jimthompson5910 (jim_thompson5910):

yep and the smallest product is -900

OpenStudy (babynini):

cool beans. Thanks!!

jimthompson5910 (jim_thompson5910):

np